You've obtained a chi square of 34.56, and it's significant. You want to test how strong the relationship between these two variables is. You have a sample size of 75. You have a table with 2 rows and 4 columns. What's your Cramer's V?
Answer #1

Formula for Cramer's V:

Here,

= 34.56

N = 75

Min(r - 1, c - 1) = Min(2 - 1, 4 - 1) = 1

Hence,

V = 0.6788